OpenStudy (anonymous):

A kindergarten school has square mats in two colors, red and blue. The ratio of the perimeter of the red mat to the perimeter of the blue mat is 2 : 1. The area of the red mat is 64 square feet. What is the area of the blue mat?

OpenStudy (tennistar):

32 square feet

OpenStudy (cwrw238):

ratio of areas of red to blue = 2^2 : 1^2 = 4 : 1 so area of blue mat = 64 4 = 16 sq ft

OpenStudy (anonymous):

The ratios are a : b and a^2 : b^2 for area. Just square the ratio and plug it in. 4 : 1 4/1 = 64/x 4x = 64 x = 16